vonnyslimedoutfr:

Two runners are saving money to attend a marathon. The first runner has $112 in savings, received a $45 gift from a friend, and will save $25 each month. The second runner has $50 in savings and will save $60 each month. Which equation can be used to find m, the number of months it will take for both accounts to have the same amount of money?

IsabelFriedmannbells:

how yo gon solve dis?

IsabelFriedmannbells:

what equation do you think you should use?

vonnyslimedoutfr:

A. 112 – 25m + 45 = 50 – 60m B. 112 + 25 + 45m = 50m + 60 C.112 + 25 – 45m = –50m + 60 D.112 + 25m + 45 = 50 + 60m

IsabelFriedmannbells:

which answer do you think it is? A. B. C. D.?

vonnyslimedoutfr:

Ice Cream idk... thats why im asking you

IsabelFriedmannbells:

well, i cant give you the answers. i mean havent you read the rules of qc?

IsabelFriedmannbells:

@aeon

IsabelFriedmannbells:

@mods

Trashy:

@isabelfriedmannbells wrote:
@aeon
hey dodo bird aeon is an ai u gotta type \ai yada yada in chat

jhonyy9:

for like a first step you need to know that this m inside equations need mean month will save $25 each month - how you think how you note this ?

IsabelFriedmannbells:

if i was yall i wuldnt help

IsabelFriedmannbells:

this person tryna get us to give them the answers

jhonyy9:

@isabelfriedmannbells wrote:
this person tryna get us to give them the answers
...direct answers note accepted ,appreciated here on QC